OpenStudy (anonymous):

find a cubic polynomial of degree 3 with real coefficients satisfying the condition f(2)=10 and whose roots include i and 3.

OpenStudy (anonymous):

*correction f(2)= -10,,,,need urgent help please

OpenStudy (blockcolder):

If one of the roots is i, then -i is also a root. Thus, our cubic equation is: \[Q(x)=a(x-i)(x+i)(x-3)\\ Q(x)=a(x^2+1)(x-3)\] Now, plug in x=2 and Q(x)=-10 to find a: \[-10=a(2^2+1)(2-3)=a(5)(-1)\\ -10=-5a \Rightarrow a=2\] Then the cubic polynomial is: \[Q(x)=2(x^2+1)(x-3)=2(x^3-3x^2+x-3)=2x^3-6x^2+2x-6\]

OpenStudy (anonymous):

thnx man

OpenStudy (blockcolder):

No problem. :D